1. One of the fundamental equations in fluid flow is the equation of discontinuity.
A. True
B. False

2. What is the equation for conservation of flow rate of an incompressible fluid through various cross sections of tube? This fluid has a speed of v_1 through a cross section A_1 and the speed of v_2 through a cross section A_2.
A. v_1A_1=v_2A_2
B. v_2A_1=v_1A_2

3. Now there is water running through a tube with two cross sections, A_1=2 meter^2 and A_2= 5 meter^2. Known that the speed v_1 at cross section A_1 is 1 meter/second, what is the value of v_2?
A. \frac{5}{2} m/s
B. \frac{2}{5} m/s

4. We have the Bernoulli equation, p_1 + \frac{\rho}{2} v_1^2=p_2+\frac{\rho}{2}v_2^2 where p_1 is the pressure when the fluid has the speed of v_1 through across section A_1 and p_2 is the pressure when the fluid has the speed of v_2 through across section A_2 , \rho is the density.

We used the equation for conservation of energy for an incompressible fluid in the tube to get Bernoulli euqation. What is the equation conservation of energy in this case? Hint: p_1V_1 is a part of energy stored in the system when the fluid has the speed of v_1 through across section A_1 and kinetic energy is the other part of this energy.(E_{k_1}=\frac{1}{2}m_1v_1^2 where m_1=\rho V_1 ).p_2V_2 is a part of energy stored in the system when the fluid has the speed of v_2 through across section A_2 and kinetic energy is the other part of this energy.(E_{k_2}=\frac{1}{2}m_2v_2^2 where m_2=\rho V_2 )

A. p_1V_1+\frac{1}{2}m_2v_2^2=p_2V_2+\frac{1}{2}m_1v_1^2
B. p_1V_1+\frac{1}{2}m_1v_1^2=p_2V_2+\frac{1}{2}m_2v_2^2

5. Now there is water running through a tube with two cross sections, A_10.2 meter^2 and A_2 = 0.4 meter^2, there is a pressure difference \Delta p_{1,2}=p_2-p_1= 15 Pa. What is value of the speed v_2 at cross section A_2? Water density is 1000 kg/meters^3. Hint: you can Bernoulli equation, p_1 + \frac{\rho}{2} v_1^2=p_2+\frac{\rho}{2}v_2^2 .

A. \sqrt{10} m/s
B. 0.01 m/s
C. 1 m/s
D. 0.1 m/s
Answer #1

1. One of the fundamental equations in fluid flow is the equation of discontinuity.
B. False

It is the equation of continuity

2.
A. v_1A_1=v_2A_2

3) As we know

v_1A_1=v_2A_2

1*2 = v2*5

v2 = 2/5 m/s


4 )

B.   p_1V_1+\frac{1}{2}m_1v_1^2=p_2V_2+\frac{1}{2}m_2v_2^2

5. Last option is correct
D. 0.1 m/s